OpenStudy (anonymous):

solve x^2=x^(1/3) show steps

OpenStudy (anonymous):

Use the log property \(loga^b = bloga\) to do it. \[ x^2=x^{(1/3)}\]\[2logx = (1/3)logx\]Then, I think you can do it.

OpenStudy (anonymous):

Still pretty lost

OpenStudy (anonymous):

Where?

OpenStudy (anonymous):

everywhere

OpenStudy (anonymous):

Do you know \[loga^b = bloga\]?

OpenStudy (anonymous):

nope

OpenStudy (mathstudent55):

\[x ^{2} = x ^{\frac{ 1 }{ 3 }}\] \[(x ^{2})^3 = (x ^{\frac{ 1 }{ 3 }})^3\]

OpenStudy (mathstudent55):

\[x ^{6} = x\] \[x ^{6} - x = 0\] \[x(x ^{5} -1) = 0\] \[x = 0 \space or \space x ^{5} =1\] x = 0 or x = 1

OpenStudy (anonymous):

Thanks

OpenStudy (mathstudent55):

wlcm